Area of a sector


Find the area of a sector of a circle having the radius of 6 cm and subtended by the central angle with measure 135°.


A circle has a radius of 10 inches and subtended by a central angle of 7π/6 radian. Find the area of a sector?


Find the area of a sector of a circle with central angle that measure 75° if the radius of a circle is 6m?


1
Expert's answer
2021-01-25T04:10:33-0500

a. Area(A) of an arc = "\pi r^2 (\frac{{1}}{360}\theta" )


A= "\pi \times 6^2 \times (\frac{{135}}{360})"


A= 42.4115 cm"^2"


b. "\theta= \frac{7}{3}\pi^r= 210^0"


area = "\pi \times 10^2 \times (\frac{{210}}{360})"


area= 183.2596 inch"^2"


c. area = "\pi \times 6^2 \times (\frac{{75}}{360})"

area = 23.5619 m"^2"